# Mastering Collaborative Multi-modal Data Selection: A Focus on Informativeness, Uniqueness, and Representativeness

### v1

@@ 4.3 Transferability of Multi-modal Data Selection
Typically, data selection for MLLMs involves selecting the most valuable data from a candidate dataset for the corresponding MLLM. The transferability analysis of multi-modal data selection aims to investigate whether the most valuable data selected for other models can be effectively transferred to the MLLM originally associated with the candidate dataset. Here, we use LLaVA-7B as the surrogate model to select valuable data from the candidate datasets of mPLUG-Owl-7B [60] and smaller Bunny-3B [17] and apply the selected data for fine-tuning these two target models to evaluate the transferability of various data selection methods. The candidate dataset of mPLUG-Owl-7B consists of 264k instances of pure text and multi-modal instruction data, while Bunny-3B’s Bunny-695k dataset contains more diverse instruction combinations. Table 2 presents the transferred results of our DataTailor and other baselines.
We observe that, despite inconsistencies between the data selection model and the target MLLMs, DataTailor still consistently achieves over 95% of the performance of the model trained on the full dataset while utilizing only 5% of the data (102.1% in mPLUG-Owl-7B and 95.6% in Bunny-3B). This demonstrates the powerful generalization capability of DataTailor and its potential in surrogate data selection. In contrast, TIVE [36] shows a significant performance drop. In contrast, TIVE performs similarly to or worse than random selection (92.0% v.s. 93.7% in Bunny-3B), although it outperforms it by a large margin in the general setting. This discrepancy may stem from TIVE’s strong correlation with training-phase gradients, making it highly sensitive to domain gaps and impairing data transferability.
@@ 4.4 In-depth Analysis
Analysis of Instruction Selection Factors. To investigate our DataTailor deeply, we study the ablation variants of different factors in Table 3. Specifically, we analyze the independence of each principle value using the following ablation strategy: 1) w/o $V_{i}^{Inf}$: we remove the informative value. 2) w/o $V_{i}^{Uni}$: we remove the unique value. 3) w/o $V_{i}^{Rep}$: we remove the representative value. 4) w/o coefficient: we remove the adaptively collaborative strategy and simply add three values with equal weight. The results of Row 3 indicate that informative value is the most crucial for multi-modal data selection. Also, Row 4 and Row 5 suggest the importance of unique values and representative values in multi-modal data selection, as unique values support MLLMs’ discriminative capabilities, while representative values enhance their generative capabilities. Furthermore, poor MLLM performance in Row 6 suggests that the adaptively collaborative strategy effectively ensures diversity across tasks in multi-modal data selection.
Methods Principled Values Benchmarks
$V_{i}^{Inf}$ $V_{i}^{Uni}$ $V_{i}^{Rep}$ MME $\uparrow$ SEED-I $\uparrow$ GQA $\uparrow$
1 Full Data - - - 1744.8 66.1 62.0
2      Random 1675.0 59.7 55.0
3      w/o $V_{i}^{Inf}$ 1712.7 60.7 57.1
4      w/o $V_{i}^{Uni}$ 1731.2 61.0 57.8
5      w/o $V_{i}^{Rep}$ 1735.5 61.4 57.9
6      w/o coefficient 1704.6 60.0 57.6
7 DataDailor 1823.7 61.7 57.7
Table 3: Ablation study of each principle in our proposed DataDailor on MLLM and VQA benchmarks. All experiments are with 7.5 % selection proportion on LLaVA-mix-665k.
@@
Robustness of Data Selection. To verify the robustness of DataTailor, we introduce two more challenging settings for multi-modal data selection: redundancy disturbance and noise disturbance. Specifically, we randomly sample 50k instructions from LLaVA-665k and construct 50k redundant data and 50k noise data through resampling and answer combination. Finally, these perturbed datasets are used as candidate datasets for data selection and the corresponding results are shown in Table 4. Empirically, DataTailor can bring out more distinctive and representative samples to identify the truly valuable samples from the redundant and noisy data for better robustness, which is crucial for discrimination tasks. Therefore, under more challenging settings with redundancy and noise disturbance, DataTailor consistently demonstrates superior performance with limited data, whereas TIVE experiences a significant performance drop (65.9 v.s. 54.1 of SciQA on redundancy disturbance and 48.5 v.s. 45.4 of GQA on noise disturbance).
Methods Redundancy Disturbance Noise Disturbance
@@
Table 4: Robustness analysis of DataTailor within redundancy disturbance and noise disturbance.
Influence of Selection Proportion kin DataTailor. As shown in Figure 5, when the selected data volume is relatively small, the model’s performance improves significantly as the data scale increases. However, due to the limited amount of valuable data in specific datasets, further increasing the data volume introduces redundancy and noise, which degrades model performance. The average performance reaches its peak at nearly 15% for LLaVA-665k and around 50% for MiniGPT4-Instruction. Moreover, we observe that on MiniGPT-4-Instruction, few samples outperform the full dataset, while performance rapidly declines when the selection ratio exceeds 50%, indicating greater data redundancy. This reveals the necessity of selecting optimal data to ensure efficiency and maintain performance.
Cross-modal Domain Clustering. Since the quality of clustering is critical for the domain-based adaptive data proportion in DataTailor, we further explore the effect of cross-modal domain clustering under different similarity thresholds of domain partition on the multi-modal data selection in Table 5. Our observations reveal that low or high similarity thresholds compromise the constraints on the uniqueness and representativeness of high-quality samples, leading to lower performance of DataTailor. Thus, we set the appropriate threshold as 0.1 for cross-modal domain clustering.
Similarity threshold w/o $V_{i}^{cor}$ 0.05 0.1 0.25 0.5
MME 1752.3 1770.3 1823.7 1782.7 1729.2
Computation Cost Analysis. Since the overhead of data selection is crucial for effective pruning methods, we analyze the advantages of DataTailor in terms of computational cost when selecting 15% data. We find that TIVE even exceeds the original training cost, which exists certain limitations. In contrast, DataTailor saves nearly 67% of the overall time while outperforming the full model’s performance. This confirms the effectiveness of our approach in reducing the instruction tuning overhead for MLLMs.
Warmup Data Selection (15%) Training
@@
Full Model - - - - $\mathcal{O}(|\mathcal{D}|\cdot|S|)$ 90 H
TIVE [36] $\mathcal{O}(|\mathcal{D}|\cdot|S_{\text{warmup}|})$ 8 H $\mathcal{O}(|\mathcal{D}|\cdot|S|)$ 100 H $\mathcal{O}(|\mathcal{D}|\cdot|S^{*}|)$ 15 H
DataTailor (Ours) - - $\mathcal{O}(|S|)$ 15 H $\mathcal{O}(|\mathcal{D}|\cdot|S^{*}|)$ 15 H
Table 6: Asymptotic complexity, wall-clock runtime (measured as 4*3090 GPU on LLaVA-665k) for total computation cost, where $|D|$ denotes the complexity of gradient computation.
Distribution of Selected Data. To give an intuitive perspective on the selected data, we employ t-SNE [50] on the feature space of selected data by DataTailor in Fig. 6. Notably, DataTailor selects informative samples without redundancy or deviation, while TIVE, despite high informativeness, focuses solely on gradient similarity, leading to redundancy and outlier noise. This visualization confirms the effectiveness of our method in selecting data that effectively adheres to three key principles.
@@ 5 Conclusion and Future Work
In this paper, we reveal the drawbacks of existing data selection methods and identify three systematic principles of informativeness, uniqueness, and representativeness as fundamental to optimizing multi-modal data selection. Building on this, we propose a unified framework, DataTailor, to synergistically integrate these principles for value evaluation and adaptively address the varying structure and complexity of samples across diverse tasks, thereby mastering collaborative multi-modal data selection. Comprehensive experiments on the challenging MLLM and general VQA benchmarks show that DataTailor significantly improves the performance of optimal data selection for MLLMs. In the future, we would like to extend DataTailor by integrating the architectural features of the surrogate model.
@@ Appendix A Overview
  • More detailed analysis of DataTailor (Section B).
@@ Appendix B DataTailor Framework
B.1 Cross-modal Domain Clustering
In this section, we introduce a cross-modal domain clustering framework designed to partition multi-modal samples into hierarchical clusters. Specifically, the framework is tailored for image-text data, which is initially grouped into different data categories based on task difficulty. Each data category is then further subdivided into multiple domains, driven by the semantic diversity within the data.
To determine the appropriate clustering for each domain, we begin by calculating the distance between different data points within that domain. We define the distance between two vectors $u$ and $v$ as:
$\text{dist}(u,v)=\sqrt{(u-v)^{2}}$ (1)
Next, we apply a hierarchical clustering algorithm to partition the data within each domain. Specifically, we utilize the Ward method to construct a spanning tree that merges cluster pairs minimizing the increase in variance. The variance increase is computed as:
$\Delta\text{SSD}=\frac{n_{A}\cdot n_{B}}{n_{A}+n_{B}}\cdot\|\overline{A}-\overline{B}\|^{2}$ (2)
where $n_{A}$ and $n_{B}$ are the sizes of the two clusters being merged, and $\overline{A}$ and $\overline{B}$ are their respective centroids. After merging clusters, the spanning tree is further divided by inspecting the sub-trees. If the maximum increase in variance $\Delta\text{SSD}$ during the merging process exceeds a predefined threshold $\lambda\Delta\text{SSD}$, the tree is split accordingly.
To optimize the construction of the spanning tree, we reduce the time complexity to $O(n^{2})$ using the nearest-neighbor chain algorithm. This approach involves adding a random starting cluster to a stack, followed by sequentially adding the closest cluster to the stack. If the two clusters at the top of the stack are the closest, they are merged and removed from the stack. This method accelerates the tree-building process while maintaining clustering accuracy.
Since the quality of clustering is critical for the domain-based adaptive data proportion in DataTailor, we investigate the impact of different values of $\lambda$ in Section 4.4, through ablation experiments to determine the most appropriate similarity threshold $\lambda$ for the cross-modal domain clustering.
B.2 Adaptive Data Proportion for Data Selection.
Since multi-modal samples exhibit varying structures and complexities across diverse tasks, we propose an adaptive weight to combine the values. Moreover, We adaptively determine the proportion of selected data for each task based on the largest singular value in spectral analysis, which empirically reflects task difficulty and gives more data selection choices to more difficult tasks in Section 3.5. In this section, we further analyze the correlation between the largest singular value and training convergence to prove that our adaptive data proportion effectively reflects task difficulty, enabling diverse data selection. We restate the previous gradient-based approach [21, 43] by employing spectral analysis to verify that the training convergence of tasks is positively correlated with the largest singular value. Assuming that the MLLM is trained using the standard cross-entropy loss and optimized with gradient descent methods, the corresponding proof is shown as follows,
$W^{{}^{\prime}}=W-\eta\nabla L(W)=W-\eta G$ (3)
$W=U\Sigma V^{\top}$ (4)
$G=U_{G}\Sigma_{G}V_{G}^{\top}$ (5)
$W^{{}^{\prime}}=U\Sigma V-\eta U_{G}\Sigma_{G}V_{G}$ (6)
$L(W^{{}^{\prime}})\approx L(W)+\langle\nabla L(W),\Delta W\rangle+\frac{1}{2}\Delta W^{\top}H\Delta W$ (7)
$L(W^{{}^{\prime}})\approx L(W)-\eta\langle\nabla L(W),G\rangle+\frac{\eta^{2}}{2}G^{\top}HG$ (8)
Assume that the softmax output is approximately linear $\|X\|=1$ with small weight changes,
$G=\frac{\partial L}{\partial W}=\sum_{i}(p_{i}-y_{i})X^{T}\approx WX^{T}$ (9)
$G\approx U\Sigma V^{\top}U_{X}\Sigma_{X}V_{X}^{\top}$ (10)
$\sigma_{max}(G)\approx\sigma_{max}(W)\sigma_{max}(X)$ (11)
Then $\sigma_{max}(G)\approx\sigma_{max}(W)$, we can use the maximum singular value of $W$ to analyze the speed of gradient descent and the change in the objective function as follows,
$L(W^{{}^{\prime}})\approx L(W)-\eta\sigma_{max}^{2}(W)+\frac{\eta^{2}}{2}\sigma_{max}^{4}(W)$ (12)
Overall, it can be seen through the gradient descent process that larger maximum singular values indicate more valuable and difficult tasks. When the singular values are larger, the norm of the gradient matrix increases, which enhances the gradient’s contribution to the objective function, thereby speeding up the gradient descent. Therefore, we assign a higher data selection proportion to more difficult tasks with larger maximum singular value ratios. Specifically, we compute the average of the largest singular value ratios for all samples in the task as follows:
$\quad x_{p}=\overline{\frac{\sigma_{0}}{\sum_{j=1}^{L_{i}}\sigma_{j}}}$ (13)

### v2

@@ 4.3 Transferability of Multi-modal Data Selection
Previous MLLM-specialized methods [43, 62] relied on the same MLLMs as in the training phase for data selection. The transferability analysis aims to investigate whether multi-modal data selected by surrogate models can be efficiently transferred into the target MLLM. Here, we use LLaVA-v1.5-7B as the surrogate model to select valuable data for the target MLLMs mPLUG-Owl-7B [67] and Bunny-3B [22], whose corresponding target datasets are mPLUG-264k and Bunny-695k, respectively. Moreover, we explore DataTailor’s robustness among other feature backbones with different structures and parameters. Table 2 presents the results of our DataTailor and other baselines
We observe that, (a) despite inconsistencies between the data selection model and the target MLLMs, DataTailor still consistently achieves over 95% relative performance of full fine-tuning with only 5% data (101.8% in mPLUG-264k and 96.0% in Bunny-695k). This demonstrates the powerful transferability of DataTailor and its potential for surrogate data selection. In contrast, TIVE [43] performs similarly to or worse than random selection (e.g., 90.0% v.s. 93.4% in Bunny-695k) though it outperforms it in Tab. 1. This discrepancy may stem from TIVE’s strong correlation with training gradients of the specific model. (b) The stable performance transition (nearly 1%) indicates that DataTailor emphasizes the inherent value of samples rather than relying on the model features as in previous feature-based baselines, which highlights its transferability for data selection.
Methods Principled Values Benchmarks
$V_{i}^{Inf}$ $V_{i}^{Uni}$ $V_{i}^{Rep}$ MME $\uparrow$ MMMU(val) $\uparrow$ SciQA $\uparrow$ Rel.
1 Full Data - - - 1744.8 32.8 70.0 100.0%
2 Random 1675.0 32.2 70.0 95.3%
3    $V_{i}^{Inf}$ 1759.3 34.9 70.2 98.0%
4    $V_{i}^{Uni}$ 1716.2 33.5 69.8 97.3%
5    $V_{i}^{Rep}$ 1771.4 33.8 68.5 97.5%
6 DataTailor 1823.7 33.9 70.9 100.1%
7    w/o adaptive collaboration 1770.2 34.0 70.2 98.8%
8    w/o adaptive proportion 1753.4 33.3 70.1 97.7%
9    w/o adaptive collaboration & proportion 1730.0 32.4 69.3 97.2%
Table 3: Ablation study of each module in DataDailor. All experiments are with 7.5 % selection proportion on LLaVA-mix-665k and Rel. denotes the average boost on all 13 benchmarks.
@@ 4.4 In-depth Analysis
Analysis of Instruction Selection Factors. To investigate our DataTailor deeply, we study the ablation variants of different factors in Table 3. Specifically, we analyze their independence using the following ablation strategy: 1) $V_{i}^{Inf}$: we only use the informative value. 2) $V_{i}^{Uni}$: we only include the unique value. 3) $V_{i}^{Rep}$: we only consider the representative value. 4) w/o adaptive collaboration & proportion: we gradually remove adaptive weights for collaboration and adaptive proportions for selection. Note that we adopt adaptive proportion in all Row 2-5 for fair comparison. The results of Row 3 indicate that informative value is the most crucial principle. Also, Row 4 and 5 suggest the importance of unique and representative values, as unique values support MLLMs’ discriminative capabilities, while representative values enhance their generative capabilities. Furthermore, decreasing performance in rows 7-9 suggests that the adaptively collaborative strategy effectively ensures diversity between tasks in multi-modal data selection.
Instantiation of Three Principles. To explore DataTailor addressing three principles for data selection, we clarify three principles based on the following insights: (a) informativeness is crucial for generalization when initializing MLLMs; (b) uniqueness is guaranteed to promote MLLMs after partial training by novel contributions; (c) representativeness improves relevant samples’ performance during training. Specifically, we instantiate three settings to evaluate each of them essentially: (a) the improvement with 1% data added to initial MLLMs reflects informativeness; (b) the improvement with extra 1% data added to MLLMs based on 15% random data training reflects uniqueness; (c) the average reduced loss of relevant samples with 1% data added reflects representativeness. We normalize each metric and compare it to three types of baselines in Fig. 1(b). The selected data from baselines show limitations in at least one principle. In contrast, DataTailor effectively selects valuable data for all three aspects of MLLM. Please refer to Appendix D.1 for more details of instantiation analysis.
@@
Robustness Analysis of DataTailor. (a) Model scale robustness. In Table 4, we verify the robustness of DataTailor with larger model scales. When keeping the same architecture and using LLaVA-v1.5-13B with higher parameters, our method demonstrates enhanced performance with larger scale MLLMs, especially +5.5 points on the LLaVA-Wild benchmark. This shows the stability of DataTailor at model scales. (b) Dataset robustness. In Table 5, we explore the dataset robustness of DataTailor with two challenging perturbations of candidate datasets: redundancy perturbation and noise perturbation. Specifically, we construct these perturbations with 50k redundant data and 50k noisy data by resampling and wrong answer combination. Note that we also sample 50k normal data from LLaVA-665k for balance. With redundancy and noise perturbations, we notice that DataTailor consistently demonstrates superior performance with limited data, whereas TIVE experiences a significant performance drop (65.9 v.s. 54.1 of SciQA on redundancy disturbance and 48.5 v.s. 45.4 of GQA on noise disturbance). It indicates DataTailor can bring out more distinctive and representative samples to identify the truly valuable samples from the redundant and noisy data for better robustness, which is crucial for discrimination tasks.
Methods MMMU(val) LLaVA-Wild SciQA GQA Rel.
LLaVA-v1.5-13B
    Full Data (100%) 35.2 69.5 72.6 63.3 100.0%
    Random  (15%) 33.0 67.3 71.8 60.7 96.3%
    DataTailor (15%) 36.4 75.0 73.9 61.2 102.4%
Table 4: Model scale robustness analysis of DataTailor.
Methods Redundancy Perturbation Noise Perturbation
@@
Table 5: Dataset robustness analysis of DataTailor with redundancy perturbation and noise perturbation.
Influence of Selection Proportion k% in DataTailor. As shown in Figure 5 (a), when the selected data volume is relatively small, the model’s performance improves as the data scale increases. However, due to the limited valuable data, further increasing the data volume introduces redundancy and noise, which degrades data-sensitive visual recognition performance (i.e. MM-Vet). This reveals the necessity of selecting data to ensure efficiency and maintain performance. Moreover, DataTailor quickly obtains over 100% overall performance (100.1% with only 7.5% data) while TIVE exhibits only limited performance while growing slowly, as shown in Figure 5 (b). It indicates that DataTailor considers the uniqueness of the selected samples to avoid repetition and achieve continuous improvement.
Computation Cost Analysis. Since the overhead of data selection is crucial for effective pruning methods, we analyze the computational cost of DataTailor when selecting 7.5% data. We find that TIVE even exceeds the original training cost, which has certain limitations. In contrast, DataTailor saves nearly 80% of the overall time while outperforming the full model’s performance. This confirms the efficiency of data selection with DataTailor for MLLMs.
Warmup Data Selection (7.5%) Training
@@
Full Model - - - - $\mathcal{O}(|\mathcal{D}|\cdot|S|)$ 100 H
TIVE [43] $\mathcal{O}(|\mathcal{D}|\cdot|S_{\text{warmup}|})$ 8 H $\mathcal{O}(|\mathcal{D}|\cdot|S|)$ 100 H $\mathcal{O}(|\mathcal{D}|\cdot|S^{*}|)$ 7.5 H
DataTailor (Ours) - - $\mathcal{O}(|S|)$ 15 H $\mathcal{O}(|\mathcal{D}|\cdot|S^{*}|)$ 7.5 H
Table 6: Asymptotic complexity and wall-clock runtime (measured with 4*3090 for LLaVA-v1.5-7B experiments on LLaVA-mix-665k dataset). $|D|$ is the complexity of gradient computation.
Distribution of Selected Data. To give an intuitive perspective on the selected data, we employ t-SNE [57] on the feature space of data chosen by DataTailor in Fig. 6. Notably, DataTailor selects informative samples without redundancy or deviation, while TIVE, despite high informativeness, focuses solely on gradient similarity, leading to redundancy and outlier noise. This visualization confirms the effectiveness of our method adhering to three principles.
@@ 5 Conclusion and Future Work
In this paper, we reveal the drawbacks of existing data selection methods and identify three systematic principles of informativeness, uniqueness, and representativeness as fundamental to optimizing multi-modal data selection. Building on this, we propose a unified framework, DataTailor, to synergistically integrate these principles for value evaluation and adaptively address the varying structure and complexity of samples across diverse tasks, thereby mastering collaborative multi-modal data selection. Comprehensive experiments on the challenging MLLM and general VQA benchmarks show that DataTailor significantly improves the performance of MLLM data selection. In the future, we aim to extend DataTailor to more challenging interleaved datasets with extra modalities such as video and audio.
Acknowledgment. This work was supported by the National Natural Science Foundation of China (62436007), the Key R&D Projects in Zhejiang Province (No. 2024C01106, 2025C01030), the Zhejiang NSF (LRG25F020001) and Wallenberg-NTU Presidential Postdoctoral Fellowship. We thank all the reviewers for their valuable comments.
@@ Appendix A Overview
  • More detailed analyses of DataTailor (Section B).
@@ Appendix B DataTailor Framework
B.1 Principle Definition
To avoid the computational infeasibility of the impact of samples on downstream task performance, we define three principles based on the geometric statistical properties of the samples to serve as practical approximation for assessing the value of multi-modal instructions:
Definition B.1 (Informativeness). It quantifies the samples’ difficulty for downstream tasks within latent space. The sample difficulty is related to its feature distribution diversity. A sample $s\in S,S\subseteq D$ is high-informativeness if its token-level feature distribution is more diverse, indicating it contains a wider variety of information at the token level. By prioritizing such samples during training, models can learn more robust representations, thereby improving inference accuracy on downstream tasks.
Intuitively, simple samples often contain redundant information (e.g., images with large areas of meaningless background or answers with repetitive descriptions). In such cases, the model can “trickily” focus on a subset of information to complete the task. Due to the high similarity of tokens in the sample, the feature matrix columns (or rows) exhibit strong linear dependence, meaning the matrix has a low rank and contains significant redundancy. According to Singular Value Decomposition (SVD), smaller singular values decrease significantly, and larger ones become disproportionately larger, leading to lower singular value entropy. Mathematically, if the feature matrix $\mathbf{M}$ has low rank, its singular value matrix $\Sigma$ can be represented as:
$\mathbf{M}=U\Sigma V$ (1)
where the smaller singular values in $\Sigma$ approach zero, while the larger singular values dominate. The singular value entropy (SVE) is defined as the entropy of normalized singular values [10], which is computed as:
$\mathbf{SVE}=-\sum_{i=1}^{r}p(\sigma_{i})\log p(\sigma_{i})$ (2)
where $\sigma_{i}$ is a singular value, $p(\sigma_{i})$ is the normalized probability distribution of the singular values, and $r$ is the rank of the singular value matrix. Since many of the singular values are close to zero, the entropy is low, reflecting the simplicity of the sample. In contrast, when the feature matrix of a sample is more information-rich and closer to full rank, singular values contribute more evenly, leading to higher entropy, which indicates a more complex sample.
Therefore, singular value entropy serves as a practical approximation of sample informativeness, enabling the selection of more challenging samples that encapsulate a diverse range of information for downstream tasks.
Definition B.2 (Uniqueness). It measures deviation from local data density. A sample $s$ is high-uniqueness in neighborhood $\mathcal{N}(s)$ if:
$\min_{s^{\prime}\in\mathcal{N}(s)}\|s-s^{\prime}\|\geq\delta\cdot\text{diam}(\mathcal{N}(s))$ (3)
where $\delta\in(0,1)$ thresholds the relative margin and $\text{diam}(\cdot)$ is the diameter of the subset within the intra-cluster space. Due to local density constraints, models are forced to learn non-degenerate decision boundaries [13], which ensures the distinction of sample distributions and improves adversarial robustness. According to Lemma 2.4 [6], maximizing the Euclidean intra-cluster margins achieves this through geometric packing in $\ell_{2}$-space, which is equivalent to sampling along the data manifold boundary to reduce redundancy between samples thereby improving training robustness [29]. It is worth noting that the dynamic computation based on the selected sample may be influenced by the greedy side effects during the selection process, making it difficult to achieve a global optimal solution.
Therefore, the static Euclidean distance between all neighborhood samples in the intra-cluster space can approximately capture data density of sample sets, allowing the selection of unique samples for continual performance improvement. Given a sample $s$ and its neighbor samples $x_{j}\in\mathbf{C}$, the Euclidean distance $d_{i,j}$ can be calculated as follows:
$d_{i,j}=\|\mathbf{p_{j}}-\mathbf{p_{i}}\|_{2}$ (4)
where a smaller Euclidean distance of two samples in the latent space indicates that these samples are highly similar, leading to a lower uniqueness value.
Definition B.3 (Representativeness). It ensures samples adhere to population-level statistics. A sample $s$ is representative if its feature vector $\phi(s)$ satisfies:
$W_{1}(\mathcal{P}_{S},\mathcal{P}_{D})\leq\gamma$ (5)
where $W_{1}$ is Wasserstein-1 distance, $\mathcal{P}_{S}$ is average feature of the subset, and $\mathcal{P}_{D}$ is average feature of the global distribution. By incorporating the Wasserstein-1 constraint, it minimizes domain shift and stabilizes the dynamics of gradient descent [7]. According to Proposition 1 [49], cosine similarity to cluster centroids approximates $W_{1}$-optimal transport under spherical normalization. Therefore, the similarity distribution from the inter-cluster space can approximately capture the alignment between samples and the overall data distribution for minimizing domain shift. The representativeness value can be measured by the cluster density of the sample, which is commonly estimated using the similarity distribution between cluster centroids:
$\tau_{i}^{c}=\frac{1}{K-1}\sum_{k\neq c}^{K}\exp({\operatorname{sim}}(\mathbf{\overline{p_{k}}},\mathbf{\overline{p_{c}}}))$ (6)
where $\mathbf{\overline{p_{c}}}$ is the feature of the cluster centroid that is calculated by the average feature of samples in the cluster. Specifically, high similarity indicates that the cluster containing the sample is well-aligned with other clusters, making it a strong representative of the overall data distribution.
B.2 Adaptive Data Proportion for Data Selection
To adapt to various task complexities within multi-modal datasets, we introduce adaptive proportion of selected data for each task. The challenge stems from the fact that task difficulty is inherently difficult to assess directly. We observe that data lacking directional diversity causes the generated trajectories to collapse into a limited subspace dominated by a few principal components. This reflects the complexity of the task, highlighting the need for more training data to enhance the robustness of MLLMs. Specifically, given the feature matrix $\mathbf{M}$ of a sample $s_{i}$, the largest singular value calculation relies on Singular Value Decomposition (SVD):
$\mathbf{M}=U\Sigma V^{T},\Sigma=\{\sigma_{1},...,\sigma_{r}\}$ (7)
where $\Sigma$ is the singular value matrix and $r$ represents the singular value matrix rank of each sample. Here $r=L_{i}$ since the sample feature dimension $d$ is much larger than the sample token length $L_{i}$ and $\sigma_{i}$ represents each singular value.
Definition B.4 (Largest Singular Value Ratio). The largest singular value ratio (LSVR) is defined as the ratio of the maximum singular value to the sum of singular values. LSVR captures the distribution of significant eigenvectors in a particular direction and reflects the task difficulty for training robustness.
$\mathbf{LSVR}=\frac{\sigma_{\text{max}}}{\sum_{j=1}^{r}\sigma_{i}}$ (8)
where $\sigma_{\text{max}}$ denotes the dominant singular value. When the LSVR becomes significantly large, it indicates insufficient variation in eigenvectors along the principal direction as follows:
$\frac{\sigma_{\text{max}}}{\sum_{i=1}^{r}\sigma_{i}}\gg\frac{1}{r}$ (9)
This indicates that information is concentrated in specific singular value directions in the latent space. Such spectral imbalance in singular values reduces the effective dimensionality of learned representations, meaning that less knowledge can be extracted from individual samples. Consequently, such difficult tasks with larger LSVR require higher data selection proportions to ensure sufficient learning. Specifically, we compute the average of the LSVR for all samples in the task as follows:
$\quad x_{p}=avg({\frac{\sigma_{max}}{\sum_{j=1}^{r}\sigma_{j}}})$ (10)